OpenStudy (anonymous):

which two solid figures have the same volume? a. a cone with a base of 5 cm2 and height of 10cm b. a rectangular solid with a base of 5cm2 and height of 10cm c. an oblique solid with a base of 5cm2 and a slant height of 10cm d. an oblique solid with a base of 5cm2 and a height of 10 cm

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you know the formulas needed?

OpenStudy (anonymous):

no i do not

jimthompson5910 (jim_thompson5910):

volume of a cone \[\Large V = \frac{1}{3}\pi*r^2*h\]

jimthompson5910 (jim_thompson5910):

Volume of Rectangular Solid \[\Large V = L*W*H\]

jimthompson5910 (jim_thompson5910):

Volume of Oblique Solid \[\Large V = A*H\] A = area of base H = height of solid

OpenStudy (anonymous):

I was thinking it was B and D

jimthompson5910 (jim_thompson5910):

what are the volumes of each?

OpenStudy (anonymous):

65.45 for the cone

OpenStudy (anonymous):

250 for rectangular solid

jimthompson5910 (jim_thompson5910):

I'm just realizing they gave you the area of the base of the cone

jimthompson5910 (jim_thompson5910):

so Area of cone = (1/3)*(area of base)*(height) Area of cone = (1/3)*(5)*(10) Area of cone = ???

OpenStudy (anonymous):

16.6

jimthompson5910 (jim_thompson5910):

yeah roughly

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

`b. a rectangular solid with a base of 5cm2 and height of 10cm` area of rectangular solid = (area of base)*(height) area of rectangular solid = (5)*(10) area of rectangular solid = ??

OpenStudy (anonymous):

50

jimthompson5910 (jim_thompson5910):

so yeah, it is B and D. Both have a volume of 50 cubic cm

jimthompson5910 (jim_thompson5910):

for some reason I have "area" on the left side when I meant to say volume

OpenStudy (anonymous):

thank you again

jimthompson5910 (jim_thompson5910):

no problem